Dan's brother is 1/2 his father's age. His brother is 22 years old. How old is his father?
Answer by stanbon(75887)

Dan's brother is 1/2 his father's age. His brother is 22 years old. How old is his father?
----
b = (1/2)f
b = 22
---
Equation:
22 = (1/2)f
father = 2*22 = 44 yrs old
==============================
Cheers,
Stan H.